OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y > -3x + 5? (0, 6) (2, 0) (-1, 9) (1, -1)

OpenStudy (anonymous):

(2,0)

OpenStudy (tamtoan):

what you need to do is to check with each pair, substitute the value of x and y in the inequality and find out which one is not right.. i believe (2,0) is not the right answer, go check again

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

plugging in each value 6 > 5 so the first one works 0 > 1 so this one doesnt work 9 > 8 so this one works -1 > 2 this on does not work

OpenStudy (tamtoan):

no, point (2,0) , sub in , you get : 0 > -3(2) + 5 = -6 + 5 = -1 so 0 > -1 ...this work also only the last one doesn't work, so the last one is the answer

OpenStudy (anonymous):

right sorry yea its the last one